# Verification Learning: Make Unsupervised Neuro-Symbolic System Feasible

### v1

@@ 6 Theoretical Study on When Knowledge Can Assist Unsupervised Learning
By minimizing the permutation error and the task-induced error, we can derive the performance bounds for knowledge validation learning in solving unsupervised neural-symbolic learning tasks.
@@
Table 1: The experiments on the dataset Addition
@@
Table 2: The time consumption in the dataset Addition.
@@
Table 3: The experiments on the dataset Chess.
Method 2 3 4 5 6
Deepproblog 49.46 33.10 25.08 49.46 15.66
NeurASP 53.66 36.07 23.93 19.07 18.82
Ground ABL 49.90 67.65 75.70 70.55 33.95
WSABL 100.00 99.80 75.95 38.30 34.35
$VL_{\perp}$ 100.00 99.90 98.00 95.70 95.15
$VL_{\perp}^{TTC}$ 100.00 99.90 92.55 91.30 92.55
$VL_{\not\perp}$ 100.00 99.90 98.00 95.70 95.15
$VL_{\not\perp}^{TTC}$ 100.00 99.90 92.55 91.30 92.55
@@ 7 Experiments
To validate the effectiveness of the framework we proposed, we conducted experiments on 4 unsupervised tasks. These experiments were primarily extensions of previously supervised Nesy tasks. For all tasks, we used LeNet as the basic network architecture (denoted as $f$) for symbol recognition from $X$ to $S$, with a learning rate of 0.001 and Adam optimizer for optimization. All experiments were completed on 4 A800 GPUs. The comparison methods include Deepproblog, Deepstochlog, NeurASP, Ground ABL, and WSABL. For algorithms that are only suitable for supervised learning, we used [True, False], i.e., whether the result matches the knowledge base, as the label for the new tasks to enable comparison. Additionally, we compared the performance of four versions of the algorithm in ablation experiments, including two Score settings under the independent and non-independent hypotheses represented by $VL_{\perp}$ and $V_{\not\perp}$, to explore the impact of optimization target settings on VL. Under the independent hypothesis, the Score setting used the neural network model’s confidence, while under the non-independent hypothesis, the Score used a combination of consistency and confidence. Furthermore, we compared the impact of using test time correction in both cases and used the tag TTC to indicate them. Our knowledge base is in the form of a verification function, and the program was written in simple Python code. During the experiments, NeurASP experienced severe timeouts (no results returned after over 300 hours), and Deepstochlog encountered memory issues (the table size limit in swi-prolog was set to $10^{12}$). In the experimental results, these issues were indicated by TLE (Time Limit Exceeded) and MLE (Memory Limit Exceeded), respectively.
@@ 7.5 Time Consumption
In addition to its excellent performance, VL also demonstrates exceptional time efficiency. We conducted a runtime comparison on the addition dataset, with time measured in seconds. The results show that VL even outperforms Ground ABL, which has preprocessing of the knowledge base. This is because VL finds the optimal solution with far fewer validation steps. The results are showed in Table 2.
@@ 8 Conclusion
We propose a knowledge verification learning framework that transforms traditional symbolic reasoning starting from supervision signal Y into a verification paradigm independent of Y. This framework is formalized as a Constraint Optimization Problem, for which we develop a dynamic combinatorial sorting algorithm. Our approach extends from scenarios satisfying independence assumptions to more general cases requiring monotonicity properties, demonstrating that monotonicity constitutes a more fundamental characteristic in Nesy. Additionally, we address the exponential growth of shortcuts through distribution alignment.
@@
The theoretical foundation establishes the problem types addressable by VL using group theory, accompanied by generalization error analysis. Experimental validation across four unsupervised learning tasks demonstrates breakthrough progress from infeasible to feasible solutions. Future work will focus on developing more sophisticated VL frameworks for complex tasks and exploring applications in broader unlabeled data scenarios.
Appendix A The Results of Additional Experiments
This section presents the experimental results that could not be included in the main text due to space limitations.
Table 4: The experiments on the dataset Sort.
Method 4 5 6 7 8
Deepproblog 96.65 97.12 95.93 39.36 45.97
DeepStochlog 84.27 81.72 77.28 MLE MLE
NeurASP 5.03 8.95 TLE TLE TLE
Ground ABL 10.20 20.22 24.19 29.73 39.35
WSABL 29.75 69.43 49.28 49.61 97.28
$VL_{\perp}$ 77.50 77.20 96.67 68.86 98.23
$VL_{\perp}^{TTC}$ 78.75 76.40 99.67 69.00 99.89
$VL_{\not\perp}$ 97.00 98.76 97.97 98.20 98.71
$VL_{\not\perp}^{TTC}$ 99.25 99.66 99.67 99.78 99.90
Table 5: The experiments on the dataset Match.
Method 6 7 8 9 10
Deepproblog 16.72 14.30 12.33 11.03 10.25
DeepStochlog 16.68 14.28 12.70 26.17 MLE
NeurASP 18.82 14.02 12.20 8.68 TLE
Ground ABL 16.68 14.28 12.35 11.03 10.02
WSABL 16.68 14.28 12.35 11.03 10.02
$VL_{\perp}$ 65.95 42.50 73.97 97.38 96.67
$VL_{\perp}^{TTC}$ 68.20 44.63 72.33 99.85 99.67
$VL_{\not\perp}$ 66.27 73.25 98.22 70.35 97.08
$VL_{\not\perp}^{TTC}$ 68.08 75.72 99.97 81.00 99.68

### v2

@@ 6.2 Data-Induced Error in VL
By minimizing the permutation error and the task-induced error, we can derive the performance bounds for VL in solving unsupervised neural-symbolic learning tasks.
@@
Table 1: The comparison of symbol recognition accuracy on the Addition dataset.
@@
Table 2: The comparison of symbol recognition accuracy on the Sort dataset.
Method 4 5 6 7 8
Deepproblog 96.65 97.12 95.93 39.36 45.97
DeepStochlog 84.27 81.72 77.28 MLE MLE
NeurASP 5.03 8.95 TLE TLE TLE
Ground ABL 10.20 20.22 24.19 29.73 39.35
A3BL 29.75 69.43 49.28 49.61 97.28
$VL_{\perp}$ 77.50 77.20 96.67 68.86 98.23
$VL_{\perp}^{TTC}$ 78.75 76.40 99.67 69.00 99.89
$VL_{\not\perp}$ 97.00 98.76 97.97 98.20 98.71
$VL_{\not\perp}^{TTC}$ 99.25 99.66 99.67 99.78 99.90
Table 3: The comparison of symbol recognition accuracy on the Match dataset.
Method 6 7 8 9 10
Deepproblog 16.72 14.30 12.33 11.03 10.25
DeepStochlog 16.68 14.28 12.70 26.17 MLE
NeurASP 18.82 14.02 12.20 8.68 TLE
Ground ABL 16.68 14.28 12.35 11.03 10.02
A3BL 16.68 14.28 12.35 11.03 10.02
$VL_{\perp}$ 65.95 42.50 73.97 97.38 96.67
$VL_{\perp}^{TTC}$ 68.20 44.63 72.33 99.85 99.67
$VL_{\not\perp}$ 66.27 73.25 98.22 70.35 97.08
$VL_{\not\perp}^{TTC}$ 68.08 75.72 99.97 81.00 99.68
Table 4: The comparison of symbol recognition accuracy on the Chess dataset.
Method 2 3 4 5 6
Deepproblog 49.46 33.10 25.08 49.46 15.66
NeurASP 53.66 36.07 23.93 19.07 18.82
Ground ABL 49.90 67.65 75.70 70.55 33.95
A3BL 100.00 99.80 75.95 38.30 34.35
$VL_{\perp}$ 100.00 99.90 98.00 95.70 95.15
$VL_{\perp}^{TTC}$ 100.00 99.90 92.55 91.30 92.55
$VL_{\not\perp}$ 100.00 99.90 98.00 95.70 95.15
$VL_{\not\perp}^{TTC}$ 100.00 99.90 92.55 91.30 92.55
Table 5: The comparison of time consumption on the Addition dataset.
@@ 7 Experiments
To validate the effectiveness of the framework we proposed, we conducted experiments on 4 unsupervised tasks. These experiments were primarily extensions of previously supervised Nesy tasks. For all tasks, we used LeNet as the basic network architecture (denoted as $f$) for symbol recognition from $X$ to $S$, with a learning rate of 0.001 and Adam optimizer for optimization. Due to the fact that many algorithms train extremely slowly, in order to conduct a performance comparison with them as comprehensively as possible, we set a unified number of epochs to 10. All experiments were completed on 4 A800 GPUs. The comparison methods include Deepproblog, Deepstochlog, NeurASP, Ground ABL, and A3BL. For algorithms that are only suitable for supervised learning, we used [True, False], i.e., whether the result matches the knowledge base, as the label for the new tasks to enable comparison. Additionally, we compared the performance of four versions of the algorithm in ablation experiments, including two Score settings under the independent and non-independent hypotheses represented by $VL_{\perp}$ and $VL_{\not\perp}$, to explore the impact of optimization target settings on VL. Under the independent hypothesis, the Score setting used the neural network model’s confidence, while under the non-independent hypothesis, the Score used a combination of consistency and confidence. Furthermore, we compared the impact of using test time correction in both cases and used the tag TTC to indicate them. Our knowledge base is in the form of a verification function, and the program was written in simple Python code. During the experiments, NeurASP experienced severe timeouts (no results returned after over 300 hours), and Deepstochlog encountered memory issues (the table size limit in swi-prolog was set to $10^{12}$). In the experimental results, these issues were indicated by TLE (Time Limit Exceeded) and MLE (Memory Limit Exceeded), respectively.
@@ 7.5 Time Consumption
In addition to its excellent performance, VL also demonstrates exceptional time efficiency. We conducted a runtime comparison on the addition dataset, with time measured in seconds. The results show that VL even outperforms Ground ABL, which has preprocessing of the knowledge base. This is because VL finds the optimal solution with far fewer verification steps. Compared with the training time of some methods that increases rapidly as the symbol space expands, the growth rate of VL’s training time is relatively slow. The results are showed in Table 5.
@@ 8 Conclusion
We propose a verification learning framework that transforms traditional symbolic reasoning starting from a supervision signal Y into a verification paradigm independent of Y. This framework is formalized as a Constraint Optimization Problem, for which we prove that it can be solved with sub-exponential complexity under the condition of independence and the more general condition of monotonicity. We propose a corresponding dynamic combinatorial sorting algorithm for this problem. We also address the exponential growth of shortcuts through distribution alignment.
@@
The theoretical foundation establishes the problem types addressable by VL using group theory, accompanied by generalization error analysis. Experimental validation across four unsupervised learning tasks demonstrates breakthrough progress from infeasible to feasible solutions.
Future work will focus on developing more sophisticated VL frameworks for complex tasks and exploring applications in broader unlabeled data scenarios.
Acknowledgements
This research was supported by the Key Program of Jiangsu Science Foundation (BK20243012), Leading-edge Technology Program of Jiangsu Science Foundation (BK20232003) and the Fundamental Research Funds for the Central Universities (022114380023).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Proof of Theorems
A.1 Proof of Theorem 4.1
We can use proof by contradiction. Suppose there exists a score-ranked $S_{i}$ at position $i$ such that $S_{i}\not\in\{S_{1},\dots,S_{i-1}\}$, and there does not exist $0<j<i$, $0<p\leq l$, $0<q<k$ such that modifying the symbol $S_{p,q}$ at position $p$ of $S_{j}$ to $S_{p,q+1}$ gives $Suc(S_{j})_{p}$ with $Score(Suc(S_{j})_{p})\geq Score(S_{i})$. Then for any $0<j<i$, there are at least two positions $p_{a}$ and $p_{b}$ where the symbols $S_{p_{a},q_{ai}}$, $S_{p_{b},q_{bi}}$ of $S_{i}$ at positions $p_{a}$ and $p_{b}$ and the symbols $S_{p_{a},q_{aj}}$, $S_{p_{b},q_{bj}}$ of $S_{j}$ at these positions satisfy $q_{ai}\geq q_{aj}+1$, $q_{bi}\geq q_{bj}+1$; or there is at least one position $p_{a}$ where the symbol $S_{p_{a},q_{ai}}$ of $S_{i}$ and $S_{p_{a},q_{aj}}$ of $S_{j}$ satisfy $q_{ai}\geq q_{aj}+2$.
For the first case, there must exist $1<j<i$ such that modifying the symbol at position $p_{a}$ (similarly for $p_{b}$) of $S_{j}$ to $S_{p_{a},q_{aj}+1}$ gives $Suc(S_{j})_{p_{a}}$. Since $q_{ai}\geq q_{aj}+1$, $q_{bi}>q_{bj}$, and for any other position $c\not\in\{a,b\}$, $q_{ci}\geq q_{cj}$, and since score is a sorted probability, we have:
$score(S_{p_{a},q_{ai}})\geq score(p_{a},q_{aj}+1),\quad score(S_{p_{b},q_{bi}})\geq score(p_{b},q_{bj}),\quad score(S_{p_{c},q_{ci}})\geq score(p_{c},q_{cj})$
By the independence assumption, $Score(S)=\prod_{i=1}^{l}score(p_{i},q_{i})$, so:
$Score(Suc(S_{j})_{p_{a}})\geq Score(S_{i})$
which leads to a contradiction, thus the assumption is invalid.
For the second case, there must exist $1<j<i$ such that modifying the symbol at position $p_{a}$ of $S_{j}$ to $S_{p_{a},q_{aj}+1}$ gives $Suc(S_{j})_{p_{a}}$. Since $q_{ai}>q_{aj}+1$ and for any other position $c\not=a$, $q_{ci}\geq q_{cj}$, we have:
$score(S_{p_{a},q_{ai}})\geq score(p_{a},q_{aj}+1),\quad score(S_{p_{c},q_{ci}})\geq score(p_{c},q_{cj})$
Thus:
$Score(Suc(S_{j})_{p_{a}})\geq Score(S_{i})$
leading to a contradiction, so the assumption is invalid. Therefore, Theorem 4.1 is proved.
A.2 Proof of Theorem 4.2
The proof process is the same as that of Theorem 4.1. Due to the satisfaction of the monotonicity assumption, for any $0<p\leq l$ and any $0<q<k$, there does not exist $S^{\prime}$ such that modifying $s_{p}$ from $s_{p,q}$ to $s_{p,q+1}$ yields $\text{Suc}(S^{\prime})_{p}=(s_{1},\dots,s_{p-1},s_{p^{\prime}},s_{p+1},\dots,s_{l})\in\text{Suc}(S^{\prime})$ satisfying $\text{Score}(\text{Suc}(S^{\prime}))>\text{Score}(S^{\prime})$. It can be seen that for any $q^{\prime}>q$, modifying $s_{p}$ from $s_{p,q}$ to $s_{p,q^{\prime}}$ to obtain $S^{\prime}_{q}$ also cannot satisfy $\text{Score}(\text{Suc}(S^{\prime}_{q}))>\text{Score}(S^{\prime})$.
Thus, for the first case, where $q_{ai}\geq q_{aj}+1$, $q_{bi}>q_{bj}$, and for any other position $c\notin\{a,b\}$, $q_{ci}\geq q_{cj}$, it can similarly be proven that $\text{Score}(\text{Suc}(S_{j})_{p_{a}})\geq\text{Score}(S_{i})$ holds, leading to a contradiction and invalidating the assumption.
For the second case, where $q_{ai}>q_{aj}+1$ and for any other position $c\neq a$, $q_{ci}\geq q_{cj}$, it can similarly be proven that $\text{Score}(\text{Suc}(S_{j})_{p_{a}})\geq\text{Score}(S_{i})$ holds, leading to a contradiction and invalidating the assumption.
Therefore, Theorem 4.2 is proved.
A.3 Proof of Theorem 4.4
The sufficiency can be proved according to Theorem 4.2.
When the monotonicity condition is not satisfied, even if we find the top-ranked $S_{1}$ by Score, we still need to calculate the scores of all $S^{\prime}\in\mathcal{S}$. It is impossible to find the maximum solution without exhaustive traversal, and even more impossible to infer the $i_{th}$ largest solution from the first $i-1$ ranked solutions. Thus, the necessity is proved.
A.4 Proof of Theorem 6.1
First, consider the clustering ability of the function $f$. We define the clustering error as the error rate between the labels assigned to samples under the optimal label assignment strategy after clustering and the true labels. According to the Rademacher complexity theory, for any function $f\in\mathcal{F}$ and a $\rho$-Lipschitz continuous loss function $L$, under the action of $n$ samples, the generalization error rate $R_{\text{oracle}}(f)$ under the optimal label assignment strategy can be determined by the empirical error rate $\hat{R}(f)$ and the model space complexity $\mathcal{R}_{n}(\mathcal{F})$, that is, with probability at least $1-\delta$, it holds that:
$R_{\text{oracle}}(f)\leq\hat{R}(f)+2\rho\mathcal{R}_{n}(\mathcal{F})+3\sqrt{\frac{\log(2/\delta)}{2n}}$
Due to the indistinguishability of symbols in the system, the error between the true label assignment strategy and the optimal label assignment strategy differs by at most $R_{\text{task}}^{\text{up}}$. The system’s generalization error $R(f)$ is determined by the generalization error $R_{\text{oracle}}(f)$ under the optimal label assignment strategy and the maximum difference $R_{\text{task}}^{\text{up}}$ between the actual label assignment strategy and the optimal one, i.e.,
$R(f)\leq R_{\text{oracle}}(f)+R_{\text{task}}^{\text{up}}$
Finally, it can be concluded that with probability at least $1-\delta$:
$R(f)\leq\hat{R}(f)+2\rho\mathcal{R}_{n}(\mathcal{F})+3\sqrt{\frac{\log(2/\delta)}{2n}}+R_{\text{task}}^{\text{up}}$